# Towards Pervasive Distributed Agentic Generative AI - A State of The Art

### v1

@@ 6. Conclusion
This paper investigated the evolving landscape of research on LLM-based agents in the domain of pervasive computing. Initially, it provided a comprehensive analysis of agents, exploring their architecture, evaluation methodologies, and diverse applications across various fields. Afterwards, the paper introduced the pervasive computing field, outlining its infrastructures and wide-ranging applications. It then explored the growing integration of artificial intelligence within this domain, culminating in the integration of intelligent agents. Building on this foundation, the paper presented a detailed examination of agent-based architectures specifically designed for pervasive environments. This included an analysis of architectural adaptations, novel strategies introduced to suit the constraints of these contexts, and a review of implemented applications. The paper then proposed a discussion of the critical challenges in this field, alongside an overview of current research efforts aimed to address these limitations.
@@
The field of artificial intelligence, is experiencing a period of exponential growth. Each month brings the release of more robust and effective models, driven by novel LLM architectures such as Mixture of Experts (Jiang et al., 2024) and increasingly efficient post-training alignment techniques such as reinforcement learning techniques like GRPO (Shao et al., 2024b), knowledge distillation, and fine-tuning methods like SFT and DPO. New powerful open-weight model families (e.g., DeepSeek r1, LLaMA 4, Qwen 3) and close-sources (e.g. Claude 4, OpenAI o1, o3) are released every few months. We are also observing the emergence of increasingly compact models that shows similar performances or even surpass larger state-of-the-art counterparts. For example, QwQ-32B demonstrates performance comparable to or better than models like DeepSeek-R1-Distilled-Qwen-32B, DeepSeek-R1-Distilled-LLaMA-70B, o1-mini, and the original DeepSeek-R1 (Qwen Team, 2025). This miniaturization is critical for pervasive deployments.
@@
Substantial progress is also being made in hardware development. Innovations like the cost-effective NVIDIA Jetson Orin and the NVIDIA DGX Spark project, featuring compact devices capable of running models with up to 200 billion parameters, clearly signal a future where affordable, personalised agent hosting is accessible to individuals and companies (NVIDIA, 2025). Also the NVIDIA CEO Jensen Huang highlights this rapid advancement, stating, ”Our systems are progressing way faster than Moore’s Law”, attributing this acceleration to integrated innovation across the entire stack of architecture, chip, systems, libraries, and algorithms (Zeff, 2025). This means that increasingly compact devices will offer greater computational power at lower costs.
@@
Over the next 2–3 years, the exponential growth in both of these areas will profoundly impact the agent field in pervasive computing, driven by the increasing performance of Fog and Edge Computing devices. The continuous integration of greater computational capabilities into smaller, more energy-efficient, and affordable hardware will enable the deployment of agents of various scales, dynamically adapting to user service requirements. Consequently, we will see a proliferation of specialized agents distributed across diverse fog and edge nodes, each providing tailored services based on their computational resources and the robustness of their embedded LLMs. This trajectory marks a definitive shift in pervasive computing: from an ”anytime, anywhere” model to one that is ”all the time, everywhere”. The traditional ”anytime, anywhere” agent approach relied on a limited number of high-performance computing (HPC) devices providing agent services, which were constrained by connectivity and response latency. In contrast, the emerging ”all the time, everywhere” agent approach leverages the increasing capability of smaller devices, such as fog and edge nodes, to host agents. This ensures the continuous provision of agent services, virtually at all times. As detailed in Section 4.1, various strategies for deploying LLMs, the core of these agents, have been proposed, as seen in works like (Shao and Li, 2025; Zhao et al., 2024b; Zhang et al., 2024f; Mitra et al., 2024), and research in this area is rapidly advancing. This future will be characterised by an increasingly interconnected network topology spanning cloud, fog, and edge devices. The enhanced performance of these edge components will enable the robust deployment of LLM-based agent services much closer to the user, facilitating faster response times and ensuring continuity of service even under unstable connectivity. Crucially, a smaller, edge-deployed agent can maintain essential functionality and deliver ongoing service in scenarios where reliance on cloud resources would be impractical or impossible. This vision lead to an increasing proliferation and usage of various agent services in multiple domains and applications, raising a crucial question:
@@
Consider a single agent attempting to manage both smartphone applications and smart home sensors. Achieving such broad expertise would demand computationally intensive, highly robust models and significant resources (a particular challenge in pervasive computing). Conversely, the demonstrated effectiveness of domain-specific fine-tuning suggests a more practical and scalable approach: deploying separate, specialised agents for distinct tasks. This lead to what we call ”Agent as a Tool”. Each specialised agent offers services that can be used like traditional software tools (Figure 5). In a realistic future scenario, each individuals can have a personal agent on their devices. These personal agents will be trained to interact with classic tools (e.g., email, calendar) (Wu et al., 2024b; Liu et al., 2024a) and, critically, call other specialised agent-tools. For instance, a personal agent might seamlessly call upon a dedicated smart home agent for managing household systems or a software engineering agent for coding assistance within a development environment. The personal agent would intelligently decide which agent-tool to use for specific subtasks. Its overall task-handling capacity would depend on factors such as the robustness of its core LLM, the computational hardware constraints (edge, fog, cloud), and the accessibility of required classic-tools or agent-tools (e.g., network availability, physical location, co-hosting status). A calendar application, for example, might be locally accessible, while specialised company agent-tools might reside on remote edge or fog devices. This vision culminates in a complex, compound system comprising personal agents, classic tools, and specialised agent-tools. While these agents will operate semi-autonomously, proposing plans that align with the specialized task’s workflow and executing corresponding actions, user interaction and guidance will remain crucial. To enhance agent performance and user experience, advanced memory strategies, such as summarizing past interactions, can be employed. This interaction data also presents a valuable opportunity for further fine-tuning or reinforcement learning techniques to improve agent performance (Wu et al., 2025). Another paramount aspect is security. For critical actions (e.g., database modifications, code commits), essential safeguards requiring authorized user approval are indispensable for both personal agents and agent-tools (Patil et al., 2024; Abdelnabi et al., 2025). Robust firewall strategies must also be in place to protect sensitive user data. Communication between personal agents, tools, and agent-tools can leverage various established protocols, such as A2A [^3] A2A, Agent Interoperability for agent-to-agent communication and the Model Context Protocol (MCP) (Anthropic, 2024) for classic tool calls. This architecture’s broad and inherent adaptability, applicable from optimizing industrial production to enhancing individual daily life, provides a significant opportunity to fundamentally reshape how businesses operate and to tangibly improve the daily experiences of people.
@@ Acknowledgements.
Funders: Fabio Ciravegna has received funding from the European Union’s HORIZON programme, project “ICOS: Towards a functional continuum operating system”, grant agreement No 101070177

### v2

@@ 7. Conclusion
This paper investigates the evolving landscape of research on LLM-based agents in the domain of pervasive computing. Initially, it provides a comprehensive analysis of agents, exploring their architecture, evaluation methodologies, and diverse applications across various fields. Afterwards, the paper introduces the pervasive computing field, outlining its infrastructures and wide-ranging applications. It then explores the growing integration of Gen-AI within this domain, culminating in the integration of intelligent agents. Building on this foundation, the paper scrutinises agent-based architectures specifically designed for pervasive environments. This includes an analysis of architectural adaptations, novel strategies introduced to suit the constraints of these contexts, and a review of implemented applications. The paper then discusses the critical challenges in this field, alongside an overview of current research efforts aimed to address these limitations.
@@
The field of artificial intelligence, is experiencing a period of exponential growth. Each month brings the release of more robust and effective models, driven by novel LLM architectures such as Mixture of Experts (jiang2024mixtral) and increasingly efficient post-training alignment techniques such as reinforcement learning techniques like GRPO (shao2024deepseekmath), knowledge distillation, and fine-tuning methods like SFT and DPO. New powerful open-weight model families (e.g., DeepSeek r1, LLaMA 4, Qwen 3) and close-sources (e.g. Claude 4, OpenAI o1, o3) are released every few months. We are also observing the emergence of increasingly compact models that shows similar performances or even surpass larger state-of-the-art counterparts. For example, QwQ-32B demonstrates performance comparable to or better than models like DeepSeek-R1-Distilled-Qwen-32B, DeepSeek-R1-Distilled-LLaMA-70B, o1-mini, and the original DeepSeek-R1 (QwenTeam_2025). This miniaturization is critical for pervasive deployments.
@@
Substantial progress is also being made in hardware development. Innovations like the cost-effective NVIDIA Jetson Orin and the NVIDIA DGX Spark project, featuring compact devices capable of running models with up to 200 billion parameters, clearly signal a future where affordable, personalised agent hosting is accessible to individuals and companies (NVIDIA_DGX_Spark_2025). Also the NVIDIA CEO Jensen Huang highlights this rapid advancement, stating, ”Our systems are progressing way faster than Moore’s Law”, attributing this acceleration to integrated innovation across the entire stack of architecture, chip, systems, libraries, and algorithms (Zeff_2025). This means that increasingly compact devices will offer greater computational power at lower costs.
@@
Over the next 2–3 years, the exponential growth in both of these areas will profoundly impact the agent field in pervasive computing, driven by the increasing performance of Edge Computing devices. The continuous integration of greater computational capabilities into smaller, more energy-efficient, and affordable hardware will enable the deployment of agents of various scales, dynamically adapting to user service requirements. Consequently, we will see a proliferation of specialized agents distributed across edge nodes, each providing tailored services based on their computational resources and the robustness of their embedded LLMs. This trajectory marks a remarkable shift in pervasive computing: from an ”anytime, anywhere” model to one that is ”all the time, everywhere”. The traditional ”anytime, anywhere” agent approach relied on a limited number of high-performance computing (HPC) devices providing agent services, which were constrained by connectivity and response latency. In contrast, the emerging ”all the time, everywhere” agent approach leverages the increasing capability of smaller devices, such as fog and edge nodes, to host agents. This ensures the continuous provision of agent services, virtually at all times. As detailed in Section 5.1, various strategies for deploying LLMs, the core of these agents, have been proposed, as seen in works like (shao2025ai; zhao2024edge; zhang2024edgeshard; mitra2024distributed), and research in this area is rapidly advancing. This future will be characterised by an increasingly interconnected network topology spanning cloud, fog, and edge devices. The enhanced performance of these edge components will enable the robust deployment of LLM-based agent services much closer to the user, facilitating faster response times and ensuring continuity of service even under unstable connectivity. Moreover, this distributed computational capacity is the key enabler for Multi-Agent Systems, where specialized agents can collaborate across the edge-to-cloud continuum to solve complex tasks that would overwhelm a single isolated model. Crucially, a smaller, edge-deployed agents can maintain essential functionality and deliver ongoing service in scenarios where reliance on cloud resources would be impractical or impossible. This vision lead to an increasing proliferation and usage of various agent services in multiple domains and applications, raising a crucial question:
@@
Consider a single agent attempting to manage both smartphone applications and smart home sensors. Achieving such broad expertise would demand computationally intensive, highly robust models and significant resources (a particular challenge in pervasive computing). Conversely, the demonstrated effectiveness of domain-specific fine-tuning suggests a more practical and scalable approach: deploying separate, specialised agents for distinct tasks. This leads to what we call ”Agent as a Tool”, where each specialized agent offers as-a-service points of provision (Figure 4). In a realistic future scenario, each individuals can have a personal agent on their devices. These personal agents will be trained to interact with classic tools (e.g., email, calendar) (wu2024avatar; liu2024apigen) and, critically, call other specialised agent-tools. For instance, a personal agent might seamlessly call upon a dedicated smart home agent for managing household systems or a software engineering agent for coding assistance within a development environment. The personal agent would intelligently decide which agent-tool to use for specific subtasks. Its overall task-handling capacity would depend on factors such as the robustness of its core LLM, the computational hardware constraints (edge, fog, cloud), and the accessibility of required classic-tools or agent-tools (e.g., network availability, physical location, co-hosting status). A calendar application, for example, might be locally accessible, while specialised company agent-tools might reside on remote edge or fog devices. This vision culminates in a complex, compound system comprising personal agents, classic tools, and specialised agent-tools. While these agents will operate semi-autonomously, proposing plans that align with the specialized task’s workflow and executing corresponding actions, user interaction and guidance will remain crucial. To enhance agent performance and user experience, advanced memory strategies, such as summarizing past interactions, can be employed. This interaction data also presents a valuable opportunity for further fine-tuning or reinforcement learning techniques to improve agent performance (wu2025collabllm). Another paramount aspect is security. For critical actions (e.g., database modifications, code commits), essential safeguards requiring authorized user approval are indispensable for both personal agents and agent-tools (patil2024goex; abdelnabi2025firewalls). Robust firewall strategies must also be in place to protect sensitive user data. Communication between personal agents, tools, and agent-tools can leverage various established protocols, such as A2A [^3] A2A, Agent Interoperability for agent-to-agent communication and the Model Context Protocol (MCP) (unknown-author-2024) for classic tool calls. This architecture’s broad and inherent adaptability, applicable from optimizing industrial production to enhancing individual daily life, provides a significant opportunity to fundamentally reshape how businesses operate and to tangibly improve the daily experiences of people.
@@ Acknowledgements.
Funders: Fabio Ciravegna has received funding from the European Union’s HORIZON programme, project “ICOS: Towards a functional continuum operating system”, grant agreement No 101070177.